# A mesh-constrained discrete point method for incompressible flows with moving boundaries

### v1

@@ 3.4 Evaluation of fluid forces acting on moving boundaries
To evaluate the surface force vector, the velocity gradients and pressure on the moving boundary $\Gamma$ are calculated using the MLS approximations. In this regard, a slight modification is made to enable compact support in the MLS reconstruction. If the symmetric 3 $\times$ 3 compact stencils are used for spatial discretization on the boundary, some meshes outside the analysis domain could be included in the compact support domain, which leads to a lack of stencils required for a stable MLS approximation (Figure 6-a). To avoid a lack of stencils, translational shifts for the compact support domain are performed as $D_{c}\rightarrow\tilde{D}_{c}$ according to a boundary normal vector $\mathbf{n}=(n_{x},~{}n_{y})$ on an interested $c$-th DP. In this study, $D_{c}$ shifts by one mesh toward the right when $n_{x}$ is positive, one mesh toward the left when $n_{x}$ is negative, and no translation occurs when $n_{x}$ is zero. The same process is followed for $n_{y}$ with top or bottom translation. Then, the interested DP is updated to $c\rightarrow c^{*}$ and the stencil-centric point $c$ is set to the centroid mesh $i$ of the parallel-shifted 3 $\times$ 3 mesh systems (Figure 6-b).
@@ 4.1 Flow past a stationary cylinder
To check the validity of the proposed method for finite inertia flows, a numerical run was conducted for flow past a stationary cylinder. The analysis domain is shown in Figure 7. The width and length of the domain were set to $H=30D$ and $L=70D$, and the distance between the center of the cylinder and the left boundary was set to $L_{\rm{c}}=20D$. The left boundary was set as an inlet ($u=U,~{}v=0$), the right boundary as an outlet ($\partial\mathbf{u}/\partial n=0,~{}P=0$), and the top and bottom boundaries and cylinder surface as a fixed wall imposing a no-slip boundary condition ($u=0,~{}v=0$). The constant inlet velocity was set to $U=1$, and the diameter of the cylinder was set to $D=1$. The dynamic viscosity $\nu$ was determined using the Reynolds number,
@@
Figure 8 shows the time changes of $C_{\rm D}$ and $C_{\rm L}$ for each Reynolds number. Good periodicity was achieved without remarkable numerical oscillation. Table 1 shows the comparisons of $C_{\rm{D}}$, $C_{\rm{L}}$, and the Strouhal number $St=fD/U$ with the vortex shedding frequency $f$. The fast Fourier transform was applied to the time histories of $C_{\rm D}$ and $C_{\rm L}$ in $t\in[100,200]$ for the calculation of the mean and standard deviation, and for frequency $f$ (using the peak frequency). Compared with the result obtained by Liu et al. [52] using a finite difference method with boundary-fitted mesh systems, the mean value of $C_{\rm D}$ is slightly overpredicted by approximately 4% for $Re=100$ and 3% for $Re=200$, and $C_{\rm L}$ was slightly different by approximately 1% for $Re=100$ and --4% for $Re=200$. Despite this, the results were in good agreement with previous results of the immersed interface method [53], the method based on the immersed boundary method [54, 55], and least-squares-based meshfree finite difference method [56]. Moreover, our results reflect the effect of the Reynolds number more appropriately than the results in [57], which are calculated by using physically consistent $\delta$-SPH method.
@@ 4.2 Single cylinder oscillation
To investigate the effect of the time resolution on fluid forces, $C_{\rm D}$ was estimated using different $\Delta t$ (Figure 13). Non-negligible numerical oscillations appeared for the pressure part in the fluid force when $\Delta t$ decreased. This is attributed to two major reasons. First, the number of updating iterations increased as $\Delta t$ decreased to reach the same physical time. This accumulated interpolation errors in the ALE formulation and broke the conservation features of the numerical solutions.
Second, the proposed method applied a collocation arrangement to the pressure projection method. This generated checkerboard oscillations for the pressure, which became remarkable when $\Delta t$ decreased. A stable formulation to suppress the oscillation [33] is required.
@@ 4.3 Two cylinders passing each other
Figure 16 shows the comparison of the time changes for the drag and lift coefficients $C_{\rm{D}}$ and $C_{\rm{L}}$, respectively, with the previous result for $t=2.5T$$3.5T$. $C_{\rm L}$ for the proposed method captured the previous result well [53]; however, a slight bias difference arose in $C_{\rm L}$ for the overall prediction.
@@ 4.5 Motion of a flapping wing
Finally, a problem was solved for complex flows with a flapping wing, initially considered in [65] to explain the mechanism of an insect hovering and developed by several researchers [65, 53, 66, 67, 54, 63]. The analysis domain is shown in Figure 22, where $c$ denotes the chord length, $e$ denotes the aspect ratio of the elliptical wing, $H$ and $L$ denote the width and length of the analysis domain, respectively, $H_{0}$ and $L_{0}$ denote the distance from the origin to the upper and right wall boundary, respectively, and $\beta$ denotes the elevation angle of the stroke plane of the elliptical wing. These parameters were set to $c=1$, $e=1$, $H=L=16c$, $H_{0}=L_{0}=6c$, and $\beta=\pi/3$. $\theta(t)$ denotes the attack angle and $a(t)$ denotes the offset distance from the origin to the rotation center of the elliptical wing:
@@
where $\theta_{0}$ is the initial attack angle, $\phi$ is the phase angle, and $a_{0}$ is the amplitude of translational motion of the wing. These parameters were set to $\theta_{0}=\pi/4$, $\phi=0$, and $a_{0}=5/2$. Note that, according to Eqs. (39) and (40), the period $T$ of the elliptical wing motion is $T=\pi a_{0}$. The width of the background mesh was set to $l_{0}=51/1600$ and the time interval $\Delta t=\pi/600$. The no-slip condition was imposed on the elliptical cylinder and the left, right upper, and lower walls. The convergence criterion for the linear system of the pressure Poisson equation was set to $1\times 10^{-3}$.
@@ 5 Conclusion
In this study, a novel meshless approach for moving boundary flow simulations in the framework of the MCD method was proposed. The proposed method represents moving boundaries by changing the role of DPs, which are constrained in background meshes. Several numerical tests were conducted to validate the proposed method, and the solutions obtained agreed well with both previous numerical and experimental results.
@@
For a future study, it is necessary to investigate reasons for numerical oscillations when $\Delta t$ is reduced. Additionally, a formulation in interaction problems between a fluid and solid needs to be applied for practical problems. It could be achieved from a numerical result that provides an accurate and less oscillating estimation of fluid forces on a moving boundary.

### v3

@@ 3.4 Evaluation of fluid forces acting on moving boundaries
To evaluate the surface force vector, the velocity gradients and pressure on the moving boundary $\Gamma$ are calculated using the MLS approximations. In this regard, a slight modification is made to enable compact support in the MLS reconstruction. If the symmetric $3\times 3$ compact stencils are used for spatial discretization on the boundary, some mesh elements outside the analysis domain could be included in the compact support domain, which leads to a lack of stencils required for a stable MLS approximation (Figure 6-a). To avoid a lack of stencils, translational shifts for the compact support domain are performed as $D_{c}\rightarrow\tilde{D}_{c}$ according to a boundary normal vector $\mathbf{n}=(n_{x},~{}n_{y})$ on an interested $c$-th DP. In this study, $D_{c}$ shifts by one mesh element toward the right when $n_{x}$ is positive, one mesh element toward the left when $n_{x}$ is negative, and no translation occurs when $n_{x}$ is zero. The same process is followed for $n_{y}$ with top or bottom translation. Then, the interested DP is updated to $c\rightarrow c^{*}$ and the stencil-centric point $c$ is set to the centroid mesh element $i$ of the parallel-shifted $3\times 3$ mesh element systems (Figure 6-b).
@@ 4 Numerical experiments
4.1 Flow past multiple stationary cylinders
A numerical example was performed for a 2D flow past stationary multiple cylinders. The main advantage of the proposed MCD method is that it reduces the computational cost of existing particle methods. The mesh constraint idea in the MCD method maintains the uniformity of the DP distribution and allows computational stencils to be compact. To investigate this feature, two types of stencils were introduced for the MLS reconstruction: the original $3\times 3$ stencil and the expanded $9\times 9$ stencil. In the latter case, the influence radius of the MLS reconstruction was set to $r_{e}=3.5l_{0}$, assuming a least-squares MPS scheme with the type-A formulation [17].
The multiple cylinders were regularly located in a square domain (Fig. 7). The inlet velocity $U=1$ was imposed at the left boundary and the outflow condition was imposed at the right boundary of the domain. The no-slip condition was imposed on the cylinder surfaces and the upper and lower walls. Here, the domain size was set to $L=1$, the width of the background mesh was set to $l_{0}=1/64$ and the time interval was set to $\Delta t=6.1\times 10^{-7}$. The number of cylinders was set to $N_{\rm cyl}=25$. Here, approximately creeping flow was considered ($Re=0.1$). The analyses were performed by changing the cylinder diameter $D_{\rm cyl}$ so that the inter-cylinder gap $h$ was between $h=12l_{0}$$2l_{0}$, and steady flow fields were obtained. The convergence criterion for the linear system of the pressure Poisson equation was set to $10^{-2}$.
Figure 7: Analysis configurations for the 2D flow past multiple stationary cylinders.
Fig. 8 shows the velocity and pressure fields using the original $3\times 3$ stencil at $h=12l_{0}$, $9l_{0}$, $6l_{0}$, and $2l_{0}$, respectively. Smooth profiles of the solution were obtained even at the smallest gap $h=2l_{0}$.
Figure 8: Velocity and pressure fields calculated by MCD method using the original $3\times 3$ stencil.
Fig. 9 shows the relationship of the porosity $\varepsilon$ and permeability $k_{\rm perm}$ defined as
$\varepsilon=1-\frac{\pi(D_{\rm cyl}/2)^{2}N_{\rm cyl}}{(2L)^{2}}=1-\frac{\pi D^{2}_{\rm cyl}N_{\rm cyl}}{16L^{2}},$ (32)
$k_{\rm perm}=\frac{2\mu UL}{\Delta P},$ (33)
where $\Delta P=p_{\rm in}-p_{\rm out}$ is the pressure difference between the inlet and outlet of the domain. The results for the original $3\times 3$ stencil and expanded $9\times 9$ stencil were nearly identical.
Figure 9: Comparison of the relationship between permeability $k_{\rm perm}$ and porosity $\varepsilon$. Green and orange circles indicate the numerical results of the MCD method using the original $3\times 3$ stencil and expanded $9\times 9$ stencil, respectively.
Table 1 shows comparisons of the number of iterations required for convergence of the pressure Poisson equation (PPE) and normalized execution time until the flow reaches steady state between the cases using the original $3\times 3$ stencil and expanded $9\times 9$ stencil for the inter-cylindrical gap $h=2l_{0}$ and $12l_{0}$. The execution time is the time elapsed from the start of the execution to a particular time step, which was set to the same number in all cases. The number of PPE iterations is given as the approximate average value for the execution. The solution with the expanded $9\times 9$ stencil reduces the number of PPE iterations by approximately half compared to the solution with the original $3\times 3$ stencil. This is believed to be the result of the wide stencil discretization communicating more information in one iteration that reduces the overall number of iterations required to converge. Meanwhile, the total execution time was inversely related, with the original $3\times 3$ stencil being roughly half that of the expanded $9\times 9$ stencil. This agrees with the fact that the number of algebraic calculations becomes small in compact stencil schemes. In general, the particle and meshless methods require wide stencils to reconstruct a polynomial that is likely to increase computational cost compared to the grid-based methods. The proposed MCD method can alleviate this problem by applying the mesh constraint idea.
Table 1: Comparisons of the number of iterations required for convergence of the pressure Poisson equation (PPE) and normalized execution time until the flow reaches steady state between the cases using the original $3\times 3$ stencil and expanded $9\times 9$ stencil for the inter-cylindrical gap $h=2l_{0}$ and $12l_{0}$. $\hat{t}$ is normalized to the execution time for the original $3\times 3$ stencils at $h=12l_{0}$.
$3\times 3$ MCD method $9\times 9$ expand stencil
$h=12l_{0}$ Iterations of PPE (approx.) 120 60
Total execution time $\hat{t}$ 1 2.63
$h=2l_{0}$ Iterations of PPE (approx.) 150 80
Total execution time $\hat{t}$ 0.697 1.16
@@ 4.2 Flow past a stationary cylinder
To check the validity of the proposed method for finite inertia flows, a numerical run was conducted for flow past a stationary cylinder. The analysis domain is shown in Figure 10. The width and length of the domain were set to $H=30D$ and $L=70D$, and the distance between the center of the cylinder and the left boundary was set to $L_{\rm{c}}=20D$. The left boundary was set as an inlet ($u=U,~{}v=0$), the right boundary as an outlet ($\partial\mathbf{u}/\partial n=0,~{}P=0$), and the top and bottom boundaries and cylinder surface as a fixed wall imposing a no-slip boundary condition ($u=0,~{}v=0$). The constant inlet velocity was set to $U=1$, and the diameter of the cylinder was set to $D=1$. The kinematic viscosity $\nu$ was determined using the Reynolds number,
@@
Figure 11 shows the time changes of $C_{\rm D}$ and $C_{\rm L}$ for each Reynolds number. Good periodicity was achieved without remarkable numerical oscillation. Table 2 shows the comparisons of $C_{\rm{D}}$, $C_{\rm{L}}$, and the Strouhal number $St=fD/U$ with the vortex shedding frequency $f$. The fast Fourier transform was applied to the time histories of $C_{\rm D}$ and $C_{\rm L}$ in $t\in[100,200]$ for the calculation of the mean and standard deviation, and for frequency $f$ (using the peak frequency). Compared with the result obtained by Liu et al. [52] using a finite difference method with boundary-fitted mesh systems, the mean value of $C_{\rm D}$ is slightly overpredicted by approximately 4% for $Re=100$ and 3% for $Re=200$, and $C_{\rm L}$ was slightly different by approximately 1% for $Re=100$ and --4% for $Re=200$. Despite this, the results were in good agreement with previous results of the immersed interface method [53], the method based on the immersed boundary method (IBM) [54, 55], and least-squares-based meshfree finite difference method [56]. Moreover, our results reflect the effect of the Reynolds number more appropriately than the results in [57], which are calculated by using physically consistent $\delta$-SPH method.
@@ 4.3 Single cylinder oscillation
The proposed MCD method involves a problem for discontinuous change of DPs due to restricting the movement of DPs, where the inner DPs assigned to describe the fluid motion suddenly appear/disappear when a moving interface crosses an edge of background mesh. This is an unavoidable feature of the MCD method that focuses on improving computational cost, and it would be a reason causing loss of numerical conservation and associated numerical dissipation and oscillation.
To investigate the effect of the time resolution on fluid forces, $C_{\rm D}$ was estimated using different $\Delta t$ (Figure 16). Non-negligible numerical oscillations appeared for the pressure part in the fluid force when $\Delta t$ decreased. The number of updating iterations increased as $\Delta t$ decreased to reach the same physical time. This accumulated interpolation errors in the ALE formulation and broke the conservation features of the numerical solutions. Moreover, the proposed method applied a collocation arrangement to the pressure projection method. This generated checkerboard oscillations for the pressure, which became remarkable when $\Delta t$ decreased. A stable formulation to suppress the oscillation [33] is required.
@@ 4.4 Two cylinders passing each other
Figure 19 shows the comparison of the time changes for the drag and lift coefficients $C_{\rm{D}}$ and $C_{\rm{L}}$, respectively, with the previous result for $t=2.5T$$3.5T$. Although a slight bias difference arose in $C_{\rm L}$ for the overall prediction, the time change behavior of $C_{\rm L}$ captured the previous result well [53].
@@ 4.6 Motion of a flapping wing
Finally, a problem was solved for complex flows with a flapping wing, initially considered in [65] to explain the mechanism of an insect hovering and developed by several researchers [65, 53, 66, 67, 54, 63]. The analysis domain is shown in Figure 25, where $c$ denotes the chord length, $e$ denotes the aspect ratio of the elliptical wing, $H$ and $L$ denote the width and length of the analysis domain, respectively, $H_{0}$ and $L_{0}$ denote the distance from the origin to the upper and right wall boundary, respectively, and $\beta$ denotes the elevation angle of the stroke plane of the elliptical wing. These parameters were set to $c=1$, $e=4$, $H=L=16c$, $H_{0}=L_{0}=6c$, and $\beta=\pi/3$. $\theta(t)$ denotes the attack angle and $a(t)$ denotes the offset distance from the origin to the rotation center of the elliptical wing:
@@
where $\theta_{0}$ is the initial attack angle, $\phi$ is the phase angle, and $a_{0}$ is the amplitude of translational motion of the wing. These parameters were set to $\theta_{0}=\pi/4$, $\phi=0$, and $a_{0}=5/2$. Note that, according to Eqs. (43) and (44), the period $T$ of the elliptical wing motion is $T=\pi a_{0}$. The width of the background mesh was set to $l_{0}=51/1600$ and the time interval $\Delta t=\pi/600$. The no-slip condition was imposed on the elliptical cylinder and the left, right upper, and lower walls. The convergence criterion for the linear system of the pressure Poisson equation was set to $2\times 10^{-3}$.
@@
The present MCD method explicitly captures moving interfaces by DPs and solves a fluid system with a common spatial discretization, i.e., MLS reconstruction. Thus, it does not require a particular treatment for spatial discretization near the interface, which is adopted in IBMs [67, 54, 63] and immersed interface method [53]. These features have an advantage in dealing with more complex and multiple boundaries. As shown in Section 4.5, a quadratic polynomial could be constructed even if there is only one fluid DP between the gap of two interfaces and this promises a stable and smooth solution. Furthermore, the common treatment of spatial discretization inside and near the interface equalizes the computational load, which is important for efficient parallel computing.
@@
4.7 An advanced algorithm for the DP arrangement
As a further development, an advanced algorithm for the DP arrangement is developed. In the original DP arrangement algorithm, the inner DPs whose masks are assigned for the inner domain are not relocated during the procedure in Eq. (9), which only relocates the boundary DPs. This procedure is updated to make the inner DPs follow the boundary surfaces as
$\mathbf{x}^{**}_{i}=\left\{\begin{array}[]{cl}\mathbf{x}^{*}_{i}+\beta\psi_{i}\hat{\mathbf{d}}_{i},&\left(i\in\Lambda^{\rm I}_{i}\right),\\[5.69054pt] \mathbf{x}^{*}_{i}+\psi_{i}\hat{\mathbf{d}}_{i},&\left(i\in\Lambda^{\Gamma}_{i}\right).\end{array}\right.$ (45)
Here, $\beta$ is the coefficient that controls the movement of DPs depending on the distance from the boundary surface, defined as
$\beta=\left\{\begin{array}[]{cl}\displaystyle\frac{\alpha}{2}\left[1+\cos{\frac{\pi\psi_{i}}{r_{e}}}\right],&{\rm if}~{}\psi_{i}<r_{e},\\[5.69054pt] 0,&{\rm otherwise},\end{array}\right.$ (46)
where the parameter $\alpha$ determines the degree of influence of $\beta$.
Fig. 28 shows the DP arrangements around a circular cylinder (or disk) at $\alpha=0$ (original DPs) and $\alpha=0.03$ (boundary-fitted DPs). Here, $r_{e}=5l_{0}$ was applied. Although both the DPs at $\alpha=0$ and 0.03 are distributed following the boundary surface because of the relaxation procedure, the boundary-fitted DPs at $\alpha=0.03$ are more smoothly distributed following the boundary shape.
Figure 28: Comparisons of the DP arrangement algorithms. (a) Original DPs ($\alpha=0$). (b) Boundary-fitted DPs ($\alpha=0.03$).
Fig. 29 shows the time change of the drag coefficient $C_{\rm D}$ for the single cylinder oscillation problem (shown in Section 4.3). In this problem, $\alpha$ was set to 0.03. Unphysical numerical oscillations were reduced using the boundary-fitted DPs. It could be considered that the smooth distribution of DPs near the boundary improves smooth interpolation using the MLS approximation.
Figure 29: Comparison of the time changes of the drag coefficient $C_{\rm D}$ for the single cylinder oscillation problem in one period (a), and their enlarged views (b, c). The orange and black lines are the results using the boundary-fitted DPs ($\alpha=0.03$) and original DPs, respectively.
Although the current approach is not the unique choice to address the boundary fitted arrangement and needs to be further verified with some numerical examples, we believe that the MCD method has a potential to handle moving boundaries more stably due to the flexibility of the DP arrangement compared to the existing Cartesian grid-based methods including IBM.
@@ 5 Conclusion
In this study, a novel meshless approach for moving boundary flow simulations in the framework of the MCD method was proposed. The proposed method represents moving boundaries by changing the role of DPs, which are constrained in the background mesh system. Several numerical tests were conducted to validate the proposed method, and the solutions obtained agreed well with both previous numerical and experimental results.
@@
  • The mesh constraint idea allows computational stencils to be compact and reduces the computational cost compared to wide-stencil schemes such as existing particle and meshless methods.
@@
An inherent limitation of the MCD method that the restriction of the DP movement by the background mesh would cause loss of numerical conservation and associated numerical dissipation and oscillation. Future research should investigate how this limitation affects practical applications such as free surface flows. Moreover, it is necessary to investigate reasons for numerical oscillations when $\Delta t$ is reduced. Additionally, a formulation in interaction problems between a fluid and solid needs to be applied for practical problems. It could be achieved from a numerical result that provides an accurate and less oscillating estimation of fluid forces on a moving boundary.